What is the power of an electric motor, in watts, if it lifts
500 g, through 80 cm in 20 seconds?
(b) What is the motor’s efficiency if the electricity supplied it
with a total of 12 J to make the lift?
1
Expert's answer
2020-12-03T10:24:25-0500

Answer

Work done in lifting mass = force × distance moved

W = weight of body × distance moved

 W= 0.5"\times9.8\times0.8"

W=3.92J

So power

"P=\frac{W}{t}=\frac{3.92}{20}=0.196 Watts"


b) efficiency

"\eta=\frac{work\space done}{suppied\space energy}=\frac{3.92}{12}"

"\eta=33\%"